Consider a two-link network where Host A is connected to the router by a 1 Mbps link with 10 ms propagation delay and the router is connected to Host B by a 15 Mbps link with a 25 ms propagation delay.

Host A sends a 1500-byte packet to Host B.

What is the transmission delay (in ms) of the packet from Host A to the router?
Answer #1

Observe: transmission delay between A and router is asked only!

a 1500 byte packet:

- over a 1Mbps link gets a transmission time = (1500 x 8)/(1 x 220) = 11.44 ms (approx)

- a propagation delay = 10 ms

Hence the total transmission delay = 11.44 + 10 = 21.44 ms (approx)